40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 16:56:38 2013, Jenne, Update, LSC, LSC model sensing matrix upgrades in progress 
    Reply  Wed Oct 9 17:27:35 2013, Jenne, Update, LSC, LSC model sensing matrix upgrades  
       Reply  Wed Oct 9 22:05:55 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
          Reply  Fri Oct 11 17:37:08 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
             Reply  Sun Oct 13 13:29:09 2013, rana, Update, LSC, New lockin / sensing matrix screens 
                Reply  Mon Oct 14 17:40:15 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
                   Reply  Thu Oct 17 13:40:55 2013, Jenne, Update, LSC, New lockin / sensing matrix frequencies All_REFL_PRMI_17Oct2013.pdf
             Reply  Mon Oct 14 21:12:35 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
          Reply  Wed Oct 16 03:04:37 2013, Jenne, Update, LSC, New lockin / sensing matrix model parts 
             Reply  Wed Oct 23 19:28:22 2013, Jenne, Update, LSC, New lockin / sensing matrix model parts 
Message ID: 9239     Entry time: Mon Oct 14 21:12:35 2013     In reply to: 9234
Author: Jenne 
Type: Update 
Category: LSC 
Subject: New lockin / sensing matrix screens 

After staring and thinking, I remembered that there is a limit to the number of characters that a channel name can have.  So, I removed the "_LOCKIN" part of the names, and recompiled, and everything seems to work.  I modified the screens that I had made, and they show all the appropriate things now.  

The symptoms were that the numbers in the filter banks (for example, INMON) were white with the usual black background.  The numbers are supposed to be green with a black background.  After I recompiled, all the numbers were green.

This also means I need to re-put in the low pass filters. 

ELOG V3.1.3-